# AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry

### v2

@@ Abstract
Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior state trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames. However, discrepancies between the prior and actual trajectory can lead to residual distortions that compromise the consistency of the LiDAR frame with its corresponding geometric environment. This imbalance may result in pointcloud registration becoming trapped in local optima, thereby exacerbating drift during long-term and large-scale localization. To address the issue, we propose a novel asymptotically and consistently converging LIO framework dubbed AC-LIO. Our key idea is to back propagate current update term based on the prior state chain, and asymptotically compensate for the residual distortion during iteration. Moreover, considering the weak correlation between previous error and current distortion, we establish convergence criteria based on the pointcloud constraints to regulate the backpropagation. This method of guiding asymptotic distortion compensation using convergence criteria subtly enhances the consistent convergence of pointcloud registration, futher improving the accuracy and robustness of LIO system. Extensive experiments demonstrate that our AC-LIO framework significantly promotes consistent convergence in state estimation compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping.
@@ I INTRODUCTION
With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of their external environments and their own motion[1, 2, 3]. As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5], humanoid robot [6] and other fields, for its high robustness and accuracy [7, 8]. Compared to Visual-Inertial Odometry, the LIO system can directly obtain depth observations of surroundings [9], without the need to estimate the depth of feature points [10, 11], and exhibits better adaptability to different lighting conditions [12, 13].
Figure 1: (a) A principal comparison of our AC-LIO with traditional approaches. Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual distortion may prevent the registration from further convergence. In contrast, our AC-LIO employs an asymptotic backpropagation to promote consistent convergence in state estimation. (b) Large scale localization and mapping result of BotanicGarden dataset. We achieve better accuracy performance than other benchmark frameworks on its sequence 1008_01 with VLP16 LiDAR.
@@
As shown in Fig. 1, traditional methods [18, 19, 20] typically only consider using IMU integration for the initial compensation of motion distortion, without further optimization of the residual distortion in the state estimation. This residual distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity. Therefore, in conventional LIO systems, such residual error may suppress further convergence of pointcloud registration during the iterative update, ultimately impacting the odometry accuracy and causing severe drift.
In this letter, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. Our AC-LIO enjoys two key technical contributions. First, during the iterative state estimation, we introduce an asymptotic distortion compensation (Sec. III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration . This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of pointcloud registration and improving the accuracy of long-term and large-scene localization.
Second, considering the weak correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec. III-D) based on the point-to-plane residuals to control the back propagation. The criteria is used to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system. Compared to other strategies such as continuous-time estimation or frame segmentation, our method leverages this criteria to construct a feedback mechanism that promotes state convergence without additional estimation, achieving a balance between localization accuracy and computational efficiency (Sec. IV-C). It can be seamlessly embedded into the iterative Kalman Filter framework [21], offering great flexibility and scalability.
@@
  • We propose a novel framework AC-LIO with asymptotic distortion compensation that showing plausible consistent convergence in long-term and large-scale scenarios.
@@
  • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios. The results show that our method achieves higher accuracy compared to current state-of-the-art frameworks.
@@ II RELATED WORK
LiDAR-Only Odometry. Studies such as LOAM [22] and F-LOAM [23] has achieved significant performance with the assumption of constant velocity motion during the corresponding time of LiDAR frames. However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth. In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence. CT-ICP [24] optimize the start and end states of LiDAR frame, and construct a continuous time trajectory by interpolation. But the real movement may not follow the interpolated trajectory, as well as the logical constraint across adjacent frames is hard to determine. I2EKF-LO [25] utilizes extra prediction iteration to estimate motion distortion, yet the update of the process noise still depends on the hyperparameters. Traj-LO [26] leverages the high frequency of LiDAR points and employs more linear segments within frame. This approach allows more accurate trajectory characterization and suppression of motion aberrations during high speed motion, but also introduces more computational cost.
@@
Continuous State LiDAR-Inertial Odometry. To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions. CLINS [33] introduces the B-spline [34] for continuous time state estimation, while SLICT [35] employs continuous-time factors to represent LiDAR constraints within a sliding window. These frameworks further optimize motion distortion with continuous-time model. However, the unconstrained trajectory would not always conform to the predefined model due to time resolution limitations in high-speed, dynamic motion. Moreover, the continuous-time optimization strategy may introduce a higher computational cost. DLIO [36] uses analytical equations to construct a continuous trajectory for precise motion correction, after coarse deskewing through IMU integration. But the method remains limited by IMU’s inherent accuracy and neglects further undistortion during the registration process.
In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation process, to achieve the asymptotic compensation of residual motion distortion. As a result, our approach can promote the consistent convergence of pointcloud registration during iterative update, further improving the accuracy over long-term and large-scale scenarios while ensuring the system robustness.
@@ III METHODOLOGY
$\mathbf{F_{a}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
@@
III-A System Overview
Figure 2: (a) The system overview of our AC-LIO framework. During iterative ESKF, asymptotic compensation of pointcloud distortion is conducted based on the convergence criteria, to promote consistent convergence of state estimation. (b) The on-manifold propagation. The blue shows the forward propagation of IMU integration. The green indicates the backward recursion of the update term according to the propagation matrix and noise relation.
III-B ESKF State Estimation
@@
$\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{a}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
@@
Considering the nonlinear errors in the $\mathbf{F_{a}}$, $\mathbf{F_{n}}$ and $\mathbf{H}$, the estimation may require iteration to promote convergence. Let $\mathbf{A}$ denote the Jacobian of update component with respect to $\delta\mathbf{x}$, and according to [18], the update equation is:
@@
With the iterative update of ESKF, our AC-LIO could further promote the consistent convergence of the state estimation via the asymptotic distortion compensation and convergence criteria introduced in following sections.
@@ III-C Asymptotic Distortion Compensation
During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise. With reference to the Rauch-Tung-Striebel fixed interval smooth strategy [39], we can back propagate the update term $\delta\mathbf{x}$ based on the prior state chain. The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$. Then, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
$\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (6)
Furthermore, this sequential backpropagation method can be extend to any states within the LiDAR frame. It allows us to selectively update several anchor in the prior state chain, thus reducing the computational complexity of backpropagation and distortion compensation. As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{a}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be:
$\begin{gathered}\mathbf{F_{a}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{a}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{a}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (7)
After each backpropagation of the update term $\hat{\mathbf{x}}_{i+n}^{j}$, the residual pointcloud distortion can continue to be compensated based on the smoothed trajectory. For the point $\mathbf{p}^{L}$ in LiDAR frame, and the state corresponding to its timestamp $t$ after the $j$th backpropagation can be denoted as $\hat{\mathbf{x}}_{t}^{j}$, then the updated coordinates can be noted as:
@@ III-D Convergence Criteria
To exploit the statistical properties of pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration. For simplification, assuming the pointcloud registration fully converges to the global optimum and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
@@
Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$. If the mean value of current point-to-plane residual is lower than this criteria, the last registration can be regarded as sufficiently converged.
Based on this threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and asymptotic distortion compensation. Noting that $\sigma_{t}=\mathbb{E}({z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
In detail, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if its APR is less than the threshold $\eta$, it can be considered to be sufficiently converged and we will continue. Then, check the convergence of current registration, if the APR is still less than $\eta$, there is no need to backpropagate for small residual distortion; otherwise, we will perform the backpropagation and compensation during the iteration.
On one hand, this approach can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-term and large-scale scenario. On the other hand, it can avoid lots of additional operations in unreliable or unnecessary conditions, to improve the system computational efficiency. The experiment results of our framework are shown in the next section.

### v3

# AC-LIO: Towards Asymptotic Compensation for Distortion in LiDAR-Inertial Odometry via Selective Intra-Frame Smoothing

@@ Abstract
Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames. However, discrepancies between the prior and true trajectory can lead to residual motion distortions that compromise the consistency of LiDAR frame with its corresponding geometric environment. This imbalance may result in pointcloud registration becoming trapped in local optima, thereby exacerbating drift during long-term and large-scale localization. To this end, we propose a novel LIO framework with selective intra-frame smoothing dubbed AC-LIO. Our core idea is to asymptotically backpropagate current update term and compensate for residual motion distortion under the guidance of convergence criteria, aiming to improve the accuracy of discrete-state LIO system with minimal computational increase. Extensive experiments demonstrate that our AC-LIO framework further enhances odometry accuracy compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping. Our code and demo are available on https://cyberkona.github.io/publication/ac-lio/.
Index Terms:
LiDAR-inertial odometry (LIO), simultaneous localization and mapping (SLAM), asymptotic distortion compensation, convergence criteria, selective intra-frame smoothing.
@@ I Introduction
With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of external environments and their own motion[1, 2, 3]. As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5] and other fields, for its high robustness and accuracy [6, 7]. Compared to Visual-Inertial Odometry, LIO system can directly obtain depth observations of surroundings, without the need to estimate the depth of feature points [8], and exhibits better adaptability to different lighting conditions [9, 10].
Figure 1: (a) A principal comparison of our AC-LIO with traditional approaches. Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual motion distortion may prevent the registration from further convergence. Our AC-LIO employs a selective intra-frame smoothing strategy to asymptotically backpropagate and compensate for residual distortion. (b) Large scale localization and mapping result of BotanicGarden dataset. We achieve better accuracy performance than other benchmark frameworks on its sequence Garden_1008 with VLP16 LiDAR.
@@
As shown in Fig. 1, traditional methods [15, 16, 17] typically only consider using IMU integration for the initial compensation of motion distortion, without further optimization of the residual distortion in the state estimation. This residual motion distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity. Therefore, in conventional LIO systems, such residual error may suppress further convergence of pointcloud registration during the iterative update, ultimately impacting the odometry accuracy and causing severe drift.
Besides, we notice that few works directly apply the reliable and concise Rauch-Tung-Striebel (RTS) smoother for continuous optimization of motion distortion within discrete state LIO frameworks. While smoothing strategies have been widely adopted in continuous state estimation and several sliding window factor graph frameworks, they primarily focus on joint optimization across multiple states, which often leads to substantial computational consumption. However, since the LIO systems pursue a balance between real-time performance and accuracy, we seek a solution tailored for discrete state LIO frameworks that, without introducing additional estimates or state variables, leverage smoothing methods to achieve asymptotical compensation for residual motion distortion during iterative state estimation. Moreover, some redundant computation can be avoid as the state covariance chains are available from IMU prior propagation.
In this work, we propose AC-LIO, a novel discrete state LIO framework with selective intra-frame smoothing. Our AC-LIO introduces two key technical contributions. First, during iterative state estimation, we propose an asymptotic distortion compensation (Sec. III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate fr the residual distortion before next iteration. This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the accuracy of long-term and large-scene localization.
Second, due to the lack of correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec. III-D) based on the point-to-plane residuals to control the back propagation. We apply the criteria to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system. Compared to other strategies such as continuous-time estimation or frame segmentation, our method leverages this criteria to construct a feedback mechanism that promotes state convergence without additional estimation, achieving a balance between localization accuracy and computational efficiency (Sec. IV-E). It can be seamlessly embedded into the iterative Kalman Filter framework, offering great flexibility and scalability.
@@
  • We propose a novel framework AC-LIO with asymptotic distortion compensation to further suppress residual distortion during iterative update, promoting system accuracy in long-term and large-scale scenarios.
@@
  • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios. The results show that our method achieves higher accuracy compared to current state-of-the-art frameworks. We also release our code to support and contribute to the open-source community.
@@ II RELATED WORK
LiDAR-Only Odometry. Studies such as LOAM [18] and F-LOAM [19] has achieved significant performance with the assumption of constant velocity motion during the corresponding time of LiDAR frames. However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth. In case of drastic velocity variation, constant velocity model cannot accurately characterize current motion state, leading to increased drift and even odometry divergence. CT-ICP [20] optimize the start and end states of LiDAR frame, and construct a continuous time trajectory by interpolation. But the real movement may not follow the interpolated trajectory, as well as the logical constraint across adjacent frames is hard to determine. I2EKF-LO [21] utilizes extra prediction iteration to estimate motion distortion, yet the update of process noise still depends on some hyperparameters. Traj-LO [22] leverages high frequency of LiDAR points and employs more linear segments within frame. This approach allows more accurate trajectory characterization and suppression of motion aberrations during high speed motion, but also introduces more computational cost.
@@
Continuous State LiDAR-Inertial Odometry. To characterize trajectories more precisely in state estimation, many frameworks come up with new optimization solutions. CLINS [29] introduces the B-spline for continuous state estimation, while SLICT [30] employs continuous-time factors to represent LiDAR constraints within a sliding window. However, the unconstrained trajectory would not always conform to predefined model due to time resolution limitations in high-speed, dynamic motion. Moreover, the continuous-time optimization strategy may introduce a higher computational cost. DLIO [31] uses analytical equations to model a continuous trajectory for precise motion correction, after coarse deskewing with IMU integration. Yet it remains limited by IMU’s inherent accuracy and neglects further undistortion during registration process.
In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation and asymptotic compensation for residual motion distortion, further improving the accuracy over long-term and large-scale scenarios while ensuring the system robustness.
@@ III METHODOLOGY
III-A System Overview
Figure 2: (a) The system overview of our AC-LIO framework. During iterative ESKF, the back propagation and asymptotic compensation for pointcloud distortion are conducted based on the convergence criteria. (b) The on-manifold propagation. The blue shows the forward propagation of IMU integration. The green indicates the backward recursion of the update term according to the propagation matrix and noise relation of error state chains.
III-B ESKF State Estimation
@@
$\mathbf{F_{p}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
@@
$\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{p}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{p}}\check{\mathbf{P}}_{i}\mathbf{F_{p}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
@@
Considering the nonlinear errors in the $\mathbf{F_{p}}$, $\mathbf{F_{n}}$ and $\mathbf{H}$, the estimation may require iteration to promote convergence. Let $\mathbf{A}$ denote the Jacobian of update component with respect to $\delta\mathbf{x}$, and according to [15], the update equation is:
@@
With the iterative update of ESKF, our AC-LIO could further promote the convergence of the state estimation via the asymptotic distortion compensation and convergence criteria introduced in following sections.
@@ III-C Asymptotic Distortion Compensation
During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise. With reference to the Rauch-Tung-Striebel fixed interval smooth strategy [34], we can back propagate the update term $\delta\mathbf{x}$ based on the prior state chain. The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{p}}\check{\mathbf{P}}_{i}\mathbf{F_{p}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$. Then, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
$\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{p}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (6)
Furthermore, this sequential backpropagation method can be extend to any states within the LiDAR frame. It allows us to selectively update several anchor in the prior state chain, thus reducing the computational complexity of backpropagation and distortion compensation. As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{p}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be:
$\begin{gathered}\mathbf{F_{p}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{p}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{p}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (7)
After each backpropagation of the update term $\hat{\mathbf{x}}_{i+n}^{j}$, where the $j$ refers the current iteration times, the residual pointcloud distortion can continue to be compensated based on the smoothed trajectory. For the point $\mathbf{p}^{L}$ in LiDAR frame, and the state corresponding to its timestamp $t$ after the $j$th backpropagation can be denoted as $\hat{\mathbf{x}}_{t}^{j}$, then the updated coordinates can be noted as:
@@ III-D Convergence Criteria
Due to the lack of correlation between the previous error and the current motion distortion within a LiDAR frame, we find it necessary to introduce the convergence criteria to regulate the asymptotic distortion compensation. If the previous LiDAR frame has sufficiently converged, the current frame may start with a relatively small initial error, allowing the backpropagation to effectively compensate for motion distortion during iterative state estimation. However, if previous LiDAR frame has not converged well, then the initial error may degrade the system accuracy via the backpropagation. Therefore, a simple and efficient method is required to detect the convergence status of each LiDAR frame, in order to guide the asymptotic distortion compensation.
@@
To exploit the statistical properties of pointcloud, during iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration. For simplification, assuming the pointcloud registration converges to the true pose and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
@@
Therefore, assuming the pointcloud registration converges to the true pose, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$. If the mean value of current point-to-plane residual is lower than this criteria, the last registration can be regarded as sufficiently converged.
Based on this threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and asymptotic distortion compensation. Noting that $\sigma_{t}=\mathbb{E}({z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice. We collect a sequence of static-movement-static behavior with Livox AVIA, to evaluate the validity of our proposed convergence criteria. As shown in Fig. 8, the value of $\sigma_{t}$ is very close to the APR during the static phases.
In detail, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if its APR is less than the threshold $\eta$, it can be considered to be sufficiently converged and we will continue. Then, check the convergence of current registration, if the APR is still less than $\eta$, there is no need to backpropagate for small residual distortion; otherwise, we will perform the backpropagation and compensation during the iteration. On one hand, this approach can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-term and large-scale scenario. On the other hand, it can avoid lots of additional operations in unreliable or unnecessary conditions, to improve the system computational efficiency. The experiment results of our framework are shown in the next section.